# FinEval-KR: A Financial Domain Evaluation Framework for Large Language Models’ Knowledge and Reasoning

### v1

@@ Appendix F Details of Experiment
We select several popular LLMs, including DeepSeek-R1, DeepSeek-V3, Qwen2.5-72B_Instruct, and QwQ-32B-preview. To study the effect of model size on performance, we also include smaller models, such as GLM-4-9B, Mistral-small-22B, and Qwen2.5-14B_Instruct. Furthermore, we evaluate Qwen2.5-14B-R1-distilled to investigate the impact of specific reasoning optimization on evaluation metrics. In total, this study includes eight open-source LLMs.
@@
We include a selection of prominent models. This selection covers the latest reasoning models, such as Claude-3.7-sonnet, Gemini-2.5-flash, Gemini-2.5-pro, GPT-o3-mini, and GPT-o1-mini. We also include the current top non-reasoning models: GPT-4.1, GPT-4o, and Qwen-max-latest. Furthermore, we add some models released in 2024 and 2023, including Moonshot-v1-8k, Moonshot-v1-128k, Doubao-pro-32k, Ernie-Bot-4.0, and GPT-3.5-turbo.
@@
Furthermore, we specifically select two financial reasoning LLMs for evaluation. The first is XuanYuan-finx1-preview from Duxiaoman AI-Lab, a Chinese financial dialogue and reasoning model designed specifically for the financial domain. It is also the first GPT-o1-like model in the financial industry. The second is Fin-R1, a financial reasoning LLM jointly developed by Shanghai University of Finance and Economics and StepFun Technology. This model is trained on Qwen2.5-7B-Instruct and designed for complex financial reasoning tasks, balancing high performance with low deployment cost.
@@
Qwen2.5-72B_Instruct 2024-09-19
DeepSeek-V3 (671B) 2025-03-24
DeepSeek-R1 (671B) 2025-01-20
@@
Appendix G Discussion of Experimental Results
We evaluate 23 LLMs listed in Appendix F using our proposed financial reasoning dataset. Table 5 presents the complete evaluation results across several metrics: Knowledge Score (KS), Reasoning Score (RS), Cognitive Scores ($\text{CS}_{1}$ to $\text{CS}_{5}$), and Task Accuracy (Acc). The $\text{CS}_{1}$ to $\text{CS}_{5}$ correspond to remembering, understanding, applying, analyzing, and evaluating in Bloom’s taxonomy, respectively.
Model/Metrics Acc KS RS $\text{CS}_{1}$ (remember) $\text{CS}_{2}$ (understand) $\text{CS}_{3}$ (apply) $\text{CS}_{4}$ (analyze) $\text{CS}_{5}$ (evaluate)
@@ Appendix G Discussion of Experimental Results
Doubao-pro-32k 0.424 0.805 0.557 0.508 0.515 0.229 0.557 0.553
Moonshot-v1-8k 0.479 0.840 0.613 0.571 0.577 0.289 0.612 0.610
Moonshot-v1-128k 0.453 0.837 0.593 0.550 0.557 0.269 0.557 0.589
Ernie-Bot-4.0 0.571 0.872 0.697 0.661 0.667 0.422 0.693 0.664
Qwen-max-latest 0.647 0.885 0.770 0.746 0.750 0.534 0.742 0.767
GPT-3.5-turbo 0.287 0.757 0.398 0.352 0.361 0.093 0.398 0.394
GPT-4o 0.669 0.884 0.799 0.776 0.780 0.578 0.798 0.795
GPT-4.1 0.824 0.952 0.905 0.894 0.895 0.788 0.891 0.904
@@
GPT-o1-mini 0.751 0.903 0.847 0.833 0.837 0.703 0.847 0.845
GPT-o3-mini 0.810 0.914 0.901 0.893 0.896 0.800 0.890 0.899
Gemini-2.5-pro 0.884 0.978 0.896 0.881 0.883 0.807 0.875 0.884
Gemini-2.5-flash 0.840 0.956 0.910 0.899 0.901 0.824 0.898 0.905
Claude-3.7-sonnet 0.788 0.939 0.873 0.856 0.860 0.731 0.855 0.871
@@
Xuanyuan-FinX1-preview 0.587 0.865 0.731 0.704 0.711 0.460 0.731 0.728
Fin-R1-7B 0.412 0.731 0.553 0.514 0.521 0.214 0.553 0.548
Table 5: The complete evaluation results across several metrics: Knowledge Score (KS), Reasoning Score (RS), Cognitive Scores ($\text{CS}_{1}$ to $\text{CS}_{5}$), and Task Accuracy (Acc).
G.1 Analysis of Knowledge Score
@@
  • Tier 1 (KS > 0.94) includes primarily top closed-source models (such as Gemini-2.5-pro/flash, GPT-4.1) and the strongest open-source inference model (DeepSeek-R1). These models show extremely high financial knowledge coverage.
  • Tier 2 (KS $\in$ [0.90, 0.94)) includes excellent closed-source models (such as Claude-3.7-sonnet, GPT-o3-mini, GPT-o1-mini) and large-scale open-source models (such as DeepSeek-V3). Their knowledge coverage is slightly lower than Tier 1 but remains at a high level.
  • Tier 3 (KS $\in$ [0.80, 0.90)) covers most mainstream open-source and closed-source models and has a relatively solid foundation in financial knowledge. The specialized financial model Xuanyuan-FinX1-preview is also in this range. This indicates its advantage in financial knowledge acquisition and provides good support for subsequent reasoning tasks.
  • Tier 4 (KS < 0.80) consists mainly of lightweight models (such as GLM-4-9B) or earlier models (such as GPT-3.5-turbo). Their financial knowledge coverage is relatively limited. The specialized financial model Fin-R1-7B is also in this tier. Its lower knowledge score likely results from limitations in parameter size or a lower match with the knowledge required by the test dataset.
In summary, the leading closed-source general large models and some open-source reasoning models perform best in financial knowledge coverage. Model parameter size significantly impacts KS. Current mainstream LLMs are mature in financial knowledge. While knowledge coverage no longer primarily distinguishes top models from other ones, it remains a fundamental ability for high-quality reasoning.
@@
G.2 Analysis of Reasoning Score
@@
  • Tier 1 (RS > 0.90) represents excellent reasoning ability. Except for GPT-4.1, the models in this tier are reasoning-optimized (such as DeepSeek-R1, Gemini-2.5-flash, GPT-o3-mini) and show outstanding performance in reasoning accuracy and the completeness of the logical chain.
  • Tier 2 (RS $\in$ [0.80, 0.90)) shows strong reasoning ability. This tier includes lightweight or earlier reasoning models (such as QwQ-32B-preview, GPT-o1-mini) and some non-reasoning but high-performing models (such as Gemini-2.5-pro, Claude-3.7-sonnet and DeepSeek-V3).
  • Tier 3 (RS $\in$ [0.60, 0.80)) includes models have some reasoning ability but often make more reasoning errors in complex tasks, such as GPT-4o, Qwen2.5-14B-R1-distilled, etc.
  • Tier 4 (RS < 0.60) consists primarily of early non-reasoning models (such as GPT-3.5-turboa, etc.) and lightweight non-reasoning models (such as GLM-4-9B, etc.). These models have a higher error rate during the reasoning.
In summary, explicit reasoning optimization is a key factor that influences the RS and a significant source of current performance differences among models. For example, the reasoning-optimized Qwen2.5-14B-R1-distilled score of 0.785 is significantly higher than the non-reasoning version of 0.688 in the same model family. The results highlight the need to focus future model optimization on the design and optimization of the reasoning pipeline to improve the reliability and stability of executing reasoning tasks.
@@
G.3 Analysis of Cognitive Score
The CS provides a systematic evaluation of models’ cognitive abilities based on Bloom’s Taxonomy, across five dimensions, that is remembering ($\text{CS}_{1}$), understanding ($\text{CS}_{2}$), applying ($\text{CS}_{3}$), analyzing ($\text{CS}_{4}$), and evaluating ($\text{CS}_{5}$). As Table 5 shows, CS scores generally exhibit a positive correlation with the KS and the RS.
At the lower cognitive levels ($\text{CS}_{1}$ and $\text{CS}_{2}$), most models achieve high scores, which indicates they have basic remembering and understanding abilities. However, the performance differences widen significantly for higher-order cognitive abilities such as applying, analyzing, and evaluating (i.e. $\text{CS}_{3}$ to $\text{CS}_{5}$). Therefore, our subsequent analysis focuses on these three metrics. We use the average of these higher-order cognitive scores, $\text{CS}_{\text{avg}}$, as the primary evaluation metric. The corresponding performance tiers show in Figure 19.
  • Tier 1 models ($\text{CS}_{\text{avg}}$ >0.85) perform excellently across all cognitive levels and show a distinct advantage particularly in higher-order abilities. DeepSeek-R1 (CS3=0.840, CS4=0.917, CS5=0.926) ranks first and Gemini-2.5-flash (CS3=0.824, CS4=0.898, CS5=0.905) ranks second. Except for GPT-4.1, the models in this tier are reasoning-optimized.
  • Tier 2 models ($\text{CS}_{\text{avg}}\in$ [0.75, 0.85)) also exhibit strong higher-order cognitive abilities, with scores slightly lower than Tier 1. Most excellent general LLMs fall into this range, such as Claude-3.7-sonnet, DeepSeek-V3, etc.
  • Tier 3 models ($\text{CS}_{\text{avg}}\in$ [0.59, 0.75) and CS3 $\in$ [0.4,0.6)) show a significant drop in their $\text{CS}_{3}$ scores compared to Tier 2, and generally fall below 0.60. The financial model Xuanyuan-FinX1-preview is in this tier, which indicates it still has some gap in higher-order cognitive abilities.
  • Tier 4 ($\text{CS}_{\text{avg}}$ < 0.59 and CS3 < 0.4) primarily includes lightweight non-reasoning models, such as GLM-4-9B, with a $\text{CS}_{3}$ of only 0.072. These models are clearly insufficient for higher-order cognitive tasks, where the ability to apply knowledge ($\text{CS}_{3}$) becomes a bottleneck.
In summary, top closed-source models and strong open-source reasoning models demonstrate robust performance across all cognitive levels and show a significant advantage in higher-order cognitive abilities. For most current models, performance in knowledge application ($\text{CS}_{3}$) remains a key bottleneck that limits the improvement of overall cognitive levels.
Figure 19: The average higher-order cognitive scores ($\text{CS}_{3}$ to $\text{CS}_{5}$) for the model.
G.4 Analysis of Task Accuracy
Task Accuracy (Acc) directly reflects the success rate of models in execute reasoning tasks. Only models that possess a broad knowledge base, strong reasoning ability, and good higher-order cognitive skills (especially application and analysis abilities) can achieve high accuracy in practical tasks. The tiers of models based on this metric are shown in Figure 20.
The three best-performing models are Gemini-2.5-pro (0.884), DeepSeek-R1 (0.860), and Gemini-2.5-flash (0.840). These models all exhibit strong reasoning capabilities, and their accuracy is significantly higher than that of non-reasoning models. In contrast, lightweight non-reasoning models such as GLM-4-9B (Acc=0.284) perform the worst. Among the financial LLMs, Xuanyuan-FinX1-preview (0.587) outperforms Fin-R1-7B (0.412), demonstrating the former’s relative advantage in handling domain-specific reasoning tasks.
@@ G.4 Analysis of Task Accuracy
G.5 Analysis of the Impact of Model Architecture on Evaluation Metrics
To analyze the impact of model architecture on metrics, we select a set of models using the Mixture-of-Experts (MoE) architecture and models using other architectures such as dense, respectively. Each group includes two open-source and two closed-source models. Their performance lists in Table 6. Note that official documentation on the architectures of Claude-3.7-sonnet and Gemini-2.5-pro is currently unavailable. Based on publicly available information Leanware (2025); OpenAI (2025); Erdil (2024); Howarth (2025); SigmaAI (2025), we categorize them into the Dense or other architecture group for this analysis.
Model/Metrics Acc KS RS $\text{CS}_{1}$ (remember) $\text{CS}_{2}$ (understand) $\text{CS}_{3}$ (apply) $\text{CS}_{4}$ (analyze) $\text{CS}_{5}$ (evaluate)
MoE architecture
DeepSeek-V3 (671B) 0.740 0.925 0.835 0.816 0.819 0.654 0.811 0.833
DeepSeek-R1 (671B) 0.860 0.960 0.928 0.917 0.920 0.840 0.917 0.926
GPT-4.1 0.824 0.952 0.905 0.894 0.895 0.788 0.891 0.904
Gemini-2.5-flash 0.840 0.956 0.910 0.899 0.901 0.824 0.898 0.905
Dense or other architecture
Qwen2.5-72B_Instruct 0.641 0.886 0.769 0.743 0.750 0.525 0.769 0.749
QwQ-32B-preview 0.731 0.901 0.850 0.832 0.838 0.674 0.850 0.847
GPT-o1-mini 0.751 0.903 0.847 0.833 0.837 0.703 0.847 0.845
GPT-o3-mini 0.810 0.914 0.901 0.893 0.896 0.800 0.890 0.899
Table 6: The performance of models for the MoE architecture and Dense or other architectures.
Effects on KS
As shown in Table 6, all of the highest KS (Knowledge Score) results are achieved by models with MoE (Mixture-of-Experts) architectures. While one reason for this might be that MoE models are generally newer, this also confirms the principle that the MoE architecture helps expand a model’s knowledge capacity by activating different expert networks that cover a broader range of knowledge domains. At the same time, compared to Dense architectures, which activate all parameters, MoE only activates a subset of experts during inference, leading to efficient knowledge storage and retrieval.
Effects on RS
In terms of reasoning ability, the MoE architecture is particularly outstanding, it occupies the top three positions. This phenomenon indicates that the MoE architecture appears more suitable for complex reasoning tasks. It suggests that different experts can specialize in specific types of reasoning patterns, and these specialized skills are combined through a dynamic routing mechanism, which enhances the accuracy and robustness of the reasoning chain.
Effects on CS
MoE models show a significant advantage across all cognitive levels, particularly in higher-order abilities such as $\text{CS}_{3}$ to $\text{CS}_{5}$. DeepSeek-R1 and Gemini-2.5-flash nearly dominate the highest scores in these dimensions. This further supports the hypothesis that the MoE architecture handles complex cognitive tasks through expert specialization. May be different types of cognitive operations (e.g., applying knowledge, analyzing reasoning processes, evaluating arguments) are performed by different experts networks.
Effects on Accuray
As a comprehensive metric of knowledge, reasoning, and cognitive abilities, MoE models naturally hold a leading position in task accuracy due to their excellent performance in the aforementioned dimensions. Their multi-expert collaboration mechanism effectively enhances the models’ ability to complete tasks in real-world scenarios.
Appendix H License and Usage Constraints
@@
Appendix I AI Assistants Usage Disclosure

### v3

@@ Appendix H Details of Experiment
We select several popular LLMs, including DeepSeek-R1, DeepSeek-V3, and QwQ-32B-preview. To study the effect of model size on performance, we also include smaller models, such as Qwen2.5-14B_Instruct. In total, this study includes four open-source LLMs.
@@
We include a selection of prominent models. This selection covers the latest reasoning models, such as Claude-3.7-sonnet, Gemini-2.5-flash, Gemini-2.5-pro, o3-mini, and o1-mini. We also include the current top non-reasoning models: GPT-4.1, GPT-4o, and Qwen-max-latest. Furthermore, we add some models released in 2024 and 2023, including Moonshot-v1-128k, Doubao-pro-32k, Ernie-Bot-4.0, and GPT-3.5-turbo.
@@
Furthermore, we specifically select two financial reasoning LLMs for evaluation. The first is Xuanyuan-FinX1-preview from Duxiaoman AI-Lab, a Chinese financial dialogue and reasoning model designed specifically for the financial domain. It is also the first o1-like model in the financial industry. The second is Fin-R1, a financial reasoning LLM jointly developed by Shanghai University of Finance and Economics and StepFun Technology. This model is trained on Qwen2.5-7B_Instruct and designed for complex financial reasoning tasks, balancing high performance with low deployment cost.
@@
DeepSeek-V3 2025-03-24
DeepSeek-R1 2025-01-20
@@
Appendix I Discussion of Experimental Results
We evaluate 18 LLMs listed in Appendix H using our proposed financial reasoning dataset. Table 7 presents the complete evaluation results across several metrics: Knowledge Score (KS), Reasoning Score (RS), Cognitive Scores (CS1 to CS5), and Task Accuracy (Acc). Results in the table are from three runs of each model. The CS1 to CS5 correspond to remembering, understanding, applying, analyzing, and evaluating in Bloom’s taxonomy, respectively.
For all subsequent analyses, our focus is on performance tiers instead of absolute scores, which helps alleviate assessment errors caused by the systematic bias and randomness of the judge model. Additionally, we set the distribution of models across the tiers to 3:4:5:6.
[…]
CS1
[…]
CS2
[…]
CS3
[…]
CS4
[…]
CS5
[…]
@@ Appendix I Discussion of Experimental Results
[…]
@@
[…]
@@
[…]
Table 7: The complete evaluation results across several metrics: Knowledge Score (KS), Reasoning Score (RS), Cognitive Scores (CS1 to CS5), and Task Accuracy (Acc).
I.1 Analysis of Knowledge Score
@@
Tier 1 is exclusively composed of closed-source models that exhibit exceptionally high financial knowledge coverage. Tier 2 is dominated by the top-performing open-source models. Although they rank just below Tier 1, the absolute score difference is marginal, indicating that their financial knowledge coverage is nearly on par with the leading closed-source models.
A significant performance drop-off occurs in the lower tiers. In Tiers 3 and 4, the older GPT-3.5-turbo notably outperforms other models within this bracket. At the bottom of the ranking is the specialized financial model, Fin-R1-7B, whose lower performance is primarily attributed to its significantly smaller parameter scale.
In summary, leading closed-source and top open-source reasoning models demonstrate the strongest performance in financial knowledge coverage, which is significantly influenced by model scale. While financial knowledge is a mature capability in most mainstream LLMs and no longer the primary differentiator among top models, it remains a fundamental prerequisite for high-quality reasoning.
@@
I.2 Analysis of Reasoning Score
@@
Tier 1 represents the pinnacle of performance, comprising reasoning-optimized models that demonstrate outstanding accuracy and logical completeness. Tier 2 includes some high-performing, non-reasoning models like GPT-4.1 and DeepSeek-V3. A significant performance gap separates the top two tiers from the bottom two. This clear stratification underscores the need for future model development to prioritize the design and optimization of the reasoning pipeline, which is crucial for enhancing the reliability and stability of complex reasoning tasks.
@@
I.3 Analysis of Cognitive Score
The CS provides a systematic evaluation of models’ cognitive abilities based on Bloom’s Taxonomy, across five dimensions, that is remembering (CS1), understanding (CS2), applying (CS3), analyzing (CS4), and evaluating (CS5). As Table 7 shows, CS scores generally exhibit a positive correlation with the KS and the RS.
While most models achieve high scores at lower cognitive levels (CS1: Remembering, CS2: Understanding), their performance diverges significantly on higher-order tasks. These more demanding abilities—Applying (CS3), Analyzing (CS4), and Evaluating (CS5)—reveal clear distinctions among the models. Consequently, our analysis focuses on these three dimensions. We define a primary metric, $\text{CS}{{}_{\text{avg}}}$, as the average score across these higher-order skills, and stratify the models into four performance tiers based on this metric (see Figure 19).
Tier 1 models excel across all cognitive levels, demonstrating a distinct advantage in higher-order abilities. This tier is led by DeepSeek-R1, followed by Gemini-2.5-flash and pro. Tier 2 models also exhibit strong higher-order cognitive skills, with performance slightly below that of Tier 1. This tier includes most general-purpose reasoning models as well as the top-performing non-reasoning models, DeepSeek-V3 and GPT-4.1. Tiers 3 and 4 primarily consist of non-reasoning or smaller-scale models.
Figure 19: The average higher-order cognitive scores (CS3 to CS5) for the model.
I.4 Analysis of Task Accuracy
Task Accuracy measures a model’s direct success rate in executing reasoning tasks. Achieving high accuracy requires a synthesis of a broad knowledge base, robust reasoning capabilities, and advanced cognitive skills—particularly in application and analysis. Consequently, the performance gradient observed in Task Accuracy closely mirrors those of the RS and CS. The tiers of models based on this metric are shown in Figure 20.
@@ I.4 Analysis of Task Accuracy
I.5 Variance Analysis
We evaluate a model’s performance stability by the standard deviation of its scores across multiple test runs. We classify stability into two categories: High Stability (a standard deviation on the order of $10^{-4}$ to $10^{-3}$), indicating highly consistent and reproducible outputs, and Low Stability (an order of $10^{-2}$ to $10^{-1}$), which suggests significant performance fluctuations.
Our key findings are as follows:
  • Knowledge Retrieval is More Stable Than Reasoning. For most models, the KS is consistently more stable than the RS. This is intuitive, as retrieving a stored fact is a more deterministic process for a well-trained model than performing a complex, multi-step logical deduction, which allows for greater variability.
  • GPT-4o is a Unique Outlier. GPT-4o defies the general trend. Its reasoning process is remarkably stable, with an RS standard deviation of $8\times 10^{-3}$, which is significantly more stable than its knowledge retrieval (KS standard deviation of $2\times 10^{-2}$). We hypothesize that GPT-4o may possess a highly consistent, almost programmatic reasoning structure, while its knowledge function exhibits greater variance to adapt to diverse queries. This unusual stability profile warrants further investigation.
Appendix J Details of Human Evaluators and Validation Process
J.1 Evaluator Qualifications and Number
A total of 30 human experts participated in our validation effort. All experts are postgraduate students with academic backgrounds in finance, economics, or statistics, ensuring they possess an accurate understanding of the relevant professional terminology, fundamental concepts, and practical scenarios.
The entire validation process was conducted on a professional annotation platform provided by a leading technology company to ensure procedural standardization and data security.
J.2 Quality Control Mechanism
To guarantee the reliability of our validation results, we implemented a multi-stage quality control process. First, we randomly sampled 10% of the dataset. Each sample was independently validated by 2 experts to ensure consistency. Following this cross-validation, we organized a team of 3 senior experts to conduct a final quality check on a random 10% of the already-validated sample (amounting to a final check on 1% of the total dataset). This final step was designed to ensure the quality and uniformity of the standards applied during the cross-validation stage.
J.3 The Validation Process
The experts’ validation work was divided into three strict, sequential stages:
Stage 1: Question and Knowledge Point Validation
In this initial stage, experts were only shown the question and its associated knowledge points. They were tasked with the following checks:
  • Question Validity: Is the question relevant to a realistic financial scenario? Is professional terminology used correctly? Are the numerical values within a reasonable range? Does the question have a single, definitive answer?
  • Knowledge Point Relevance: Are the tagged knowledge points accurate and comprehensive? Is the naming of the knowledge points consistent with standard terminology in mainstream textbooks?
Independent Answering
After confirming the quality of the question, experts were required to solve the problem independently, without reference to any provided solution. The goal of this step was to obtain a high-quality, unbiased human answer to serve as a benchmark for subsequent comparisons.
Stage 3: Reasoning Steps and Cognitive Labels Validation
Finally, the system presented the experts with the answer, the step-by-step solution, and the Bloom’s Taxonomy cognitive label from our dataset. The experts were required to perform the following checks:
  • Answer and Solution Process Verification: First, they compared their own answer to the one in the dataset. If the answers did not match, the sample was immediately flagged as “unqualified”. If the answers matched, they proceeded to meticulously review the solution steps provided in the dataset, assessing whether the logic was clear, the steps were reasonable, and the calculations were correct.
  • Cognitive Label Accuracy Check: Based on the predefined verb list corresponding to each cognitive level (as defined in Figure 9), the experts had to judge whether the cognitive label assigned to the question was accurate.
Appendix K License and Usage Constraints
@@ I.5 Variance Analysis
Appendix L AI Assistants Usage Disclosure